# A 2-step Framework for Automated Literary Translation Evaluation: Its Promises and Pitfalls

### v2

@@ Abstract
In this work, we propose and evaluate the feasibility of a two-stage pipeline to evaluate literary machine translation, in a fine-grained manner, from English to Korean. The results show that our framework provides fine-grained, interpretable metrics suited for literary translation and obtains a higher correlation with human judgment than traditional machine translation metrics. Nonetheless, it still fails to match inter-human agreement, especially in metrics like Korean Honorifics. We also observe that LLMs tend to favor translations generated by other LLMs, and we highlight the necessity of developing more sophisticated evaluation methods to ensure accurate and culturally sensitive machine translation of literary works.
@@ 1 Introduction
While new approaches such as agentic systems are being utilized for literary translation Wu et al. (2024); He et al. (2024), evaluation remains challenging because automated metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020) are not capable of capturing the nuance of literary translation Thai et al. (2022). Recently, however, pretrained LLMs have become more common in evaluating machine translation Kocmi and Federmann (2023); Freitag et al. (2023), in lieu of and in collaboration with human evaluation.
Yet, for the agentic translation systems to work at scale, we need an automatic evaluation method first, which will give signal to the LLM agents and pipeline developer whether the translation meets the required criteria or not since human evaluation in each iteration is costly.
@@
With extensive human evaluation, focusing on English to Korean setting, we demonstrate that while the current LLMs get close to human preference in some dimensions, they still can not replace human evaluation altogether, specifically for some critical dimensions like Korean honorifics. However, we still believe our two-step literary evaluation framework will be useful for future researchers.
@@ 2 Background
Building on James S. Holmes’s 1970s’ foundational theories of literary translation, scholars like Susan Bassett, André LeFevere, and Lawrence Venuti have led the “cultural turn” in translation studies, which emphasizes the importance of cultural and historical contexts in literary translation Snell-Hornby (2006). Literary translation has a special place within broader translation studies for its focus on the aesthetic and affective use of language as well as its engagement with fictional content. Some of the persisting concerns in literary translation include the tension between literal (word-for-word) and free or adaptive (sense-for-sense) Munday et al. (2022), the visibility of the translator Venuti (2017), and the social role of the translator Blakesley (2018). More recently, scholars including Kolb and Macken have examined the potential of the integration of technology and machine translation into literary translation processes Kolb (2023); Macken et al. (2022). Our work presents a pipeline incorporating the cultural and aesthetic aspects of literary translation into the evaluation of machine translation.
@@
String-based metrics like BLEU Papineni et al. (2002) have been widely used in machine translation. Similarly, more recent metrics such as BLEURT Sellam et al. (2020), COMET Rei et al. (2020), and XCOMET Colombo et al. (2023) leverage BERT-style encoder language models Devlin et al. (2019) fine-tuned on translation preference dataset. However, previous work has shown that these metrics are not suitable for literary translation Thai et al. (2022). Departing from earlier trends, Kocmi and Federmann (2023) employed the cutting-edge LLM GPT4 OpenAI (2023) to assess machine translation, where their approach, GEMBA-MQM, set a new benchmark for translation quality estimation at WMT 2023 Freitag et al. (2023). Recent studies from Wu et al. (2024) and Chen et al. (2024) also use GPT4 as a judge for literary translation for web novels and Chinese poetry respectively.
We take a step back and conduct a comprehensive study on the LLMs’ capabilities of fine-grained evaluation for literary translation. A contemporary work, Zhang et al. (2024), has also looked into LLM’s ability to evaluate literary text.
@@
Our proposed question-answering-based translation evaluation method relates closely to automatic metrics like TIFA Hu et al. (2023), VPEval Cho et al. (2024), and DSG Cho et al. (2023) for measuring the consistency of text-to-image generation models such as DALL-E Ramesh et al. (2021) and Stable Diffusion Podell et al. (2024). In this line of work, an LLM first generates a list of verification questions from the given user’s prompt, for example “is there a car?" “is the car blue?" and then passes the questions to a vision language model such as BLIP Li et al. (2022) which answers the verification questions. This approach also has a resemblance to how previous works have measured faithfulness in summarization Kim et al. (2024); Wang et al. (2020) and long-form factuality evaluation Min et al. (2023). To the best of our knowledge, our work is the first to explore this question-answering paradigm for machine translation evaluation.
@@ 3 Dataset, Metrics & Annotators
Human Annotators:
Three literary experts with extensive experience in both English and Korean literature develop and validate the framework presented in Section 4. They also carry out the annotations in Section 5 to validate the framework. All annotators were female and native Korean speakers.
@@ 4.1 Step 1: RULER
We first conduct a small-scale qualitative evaluation of machine-translated literary texts to identify the capabilities that evaluation models should possess, with a particular focus on the types of errors they should detect. For this, we select five stories (marked dark gray in Table 7) and translate them using various large and small language models, as well as commercially available translation services, like Google Translate, to get a diverse set of translations. We observe that LLMs show inconsistent quality in doing literary translation. Larger models are more likely to produce remarkably sophisticated translations, while smaller ones often make unexpected errors, even with relatively simple passages. Our analysis identifies four major types of errors machine translation systems make when translating literary works, some of which are especially pronounced in English-to-Korean translations: word/phrase/expression-level issues, context-related issues, issues relevant to honorifics, and form-related issues (see examples in Table 1). We provide a detailed explanation of these categories with concrete examples in Appendix B.
@@
The detailed rubric for each criterion can be found in Appendix C.
@@ 4.2 Step 2: VERSE
Eliminating common errors is insufficient for achieving high-quality literary translation. A successful translation must also maintain the essential and unique literary qualities of the original work, which can vary significantly between different texts. To capture such nuances, we propose VERSE, a question-answering-based evaluation framework. We prompt a high-performing LLM (in our case GPT4-o OpenAI (2024)) to generate a list of verification questions that must be satisfied in a good translation. We instruct the question-generating model to focus on the literary aspects of the passage, with a brief summary of the story provided for context [^1] The full prompt is provided in Appendix J.1. Figure 1 provides examples of the questions generated by the LLM222additional examples are available in Appendix K. A thorough examination by native Korean speakers confirmed that LLMs are capable of generating questions that effectively address key considerations in translating literary works with varying styles and other quirks.
@@
Information about the quality of LLM generated question can be found in Appendix H.1.2.
@@
To provide better interpretability for the LLM generated questions in VERSE, we classify the questions into nine categories using an LLM. For all our experiments, the questions are classified in a zero shot manner by GPT4-o. The categories [^3] Notice that the categories are not necessarily mutually exclusive and one question might fit into more than one category., as shown in Figure 2, are defined by a human expert after qualitatively examining many evaluation questions. To validate the accuracy of the LLM classification, one human expert manually annotates 150 questions into the nine categories in which GPT4-o scores  77% Top-1 match in accuracy. An examination of the misclassified examples reveals that most misclassifications also fall into a closely related category  [^4] See confusion matrix in Figure 5.
5 Results
@@
We first create a diverse set of translations by machine translating the stories in our datasets with various LLMs and commercial translation service (e.g., Google Translate). The LLMs ranged from small 7B models to GPT4-o and Claude Sonnet 3.5555similar to what we described in Section 4.1. We also vary different prompting techniques (sentence level, paragraph level, providing extra summary, zero shot, few shot (see Appendix G)) to test the generalizability.
We then randomly select 20 ( machine translated paragraph) pairs from each of the 10 stories marked gray in Table 7 for human annotation (200 paragraphs in total). We also randomly select three evaluation questions corresponding to those paragraphs for annotation for Step 2 (600 questions in total).
Because of the subjective nature of evaluating literary translation, we create an extensive annotation guideline alongside the rubric so that annotators have the same understanding of the given task. The annotation guideline and annotation interface can be found in Appendix E. Table 2 shows that the annotators have high inter-annotator agreement. We also ask the annotators to re-annotate a subset of 40 paragraphs after four weeks of the first annotation, and the intra-annotator agreement was  0.7.
Criteria Kendall’s Spear. MSE Krip.
@@
Honorifics .70 .74 .89 .73
Lex. Choices .72 .80 .71 .79
Syntax .66 .75 .88 .74
Content .72 .81 .77 .79
LLM QnA .64 .69 .42 .69
Table 2: Inter-annotator agreement metrics across categories. Higher values indicate better agreement for Kendall’s $\tau$, Spearman’s $\rho$, and Krippendorff’s $\alpha$, while lower values are better for MSE. The values are averaged over three annotators. Notice that the last row, LLM QnA, belonging to Step 2, is measured between 1 to 3 while the others, belonging to Step 1, are measured on a 1 to 5 scale.
@@
We compare several open-weight and proprietary models, which are known for their good performance in Korean, as evaluators for our experiments. We set the temperature to 0.0 for all experiments in this section for reproducibility and report the numbers from a single run from here on.
@@
According to the common practice of literary machine translation Karpinska and Iyyer (2023); Chen et al. (2024); He et al. (2024) , we report COMET Rei et al. (2022), BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) as the baseline. However, the comparison may not be entirely appropriate because these metrics are meant to be used in sentence-level evaluation, and our dataset is aligned at paragraph-level.
@@ 5.1 RULER: Can LLMs Capture Common Mistakes in Literary Machine Translation?
Honor. 1 31 0.79 0.07 0.81 0.66
2 3 0.15 0.10
3 7 0.11 0.04
4 11 0.22 0.14
5 140 0.91 0.88
Lexical 1 16 0.54 0.43 0.54 0.47
2 32 0.47 0.45
3 45 0.45 0.37
4 47 0.50 0.52
5 52 0.70 0.51
Syntax 1 23 0.61 0.39 0.52 0.43
2 32 0.55 0.45
3 49 0.50 0.27
4 43 0.38 0.44
5 45 0.60 0.56
Content 1 19 0.57 0.19 0.54 0.46
2 32 0.41 0.47
3 47 0.45 0.22
4 48 0.47 0.43
5 46 0.73 0.70
@@
Table 3: Performance comparison across different evaluation criteria. H-F1 represents the average F1-score between human annotator pairs (3 pairs), while LLM-F1 shows the average F1-score between human annotators and GPT4o (3 pairs). H-Acc and LLM-Acc show the average accuracy for human-human and human-LLM comparisons respectively. Support (Sup.) shows the average number of instances for each label across three annotators (rounded).
Honorifics Lexical Syntax Content Aggregated Avg.
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
claude-3-5-sonnet .39 .41 2.30 .58 .65 1.17 .57 .65 1.20 .57 .64 1.38 .53 .59 1.51
claude-3-5-haiku .41 .44 1.86 .54 .60 1.37 .57 .64 1.23 .55 .62 1.42 .52 .58 1.47
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
claude-3-haiku .20 .21 2.24 .45 .50 1.57 .42 .47 1.47 .50 .56 1.55 .39 .43 1.71
gemma-2-27b-it .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
gemma-2-9b-it .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
llama-3-8b-it -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
COMET .09 .11 .27 .34 .24 .32 .27 .35 .22 .28
BERTScore .12 .15 .23 .29 .22 .28 .24 .31 .20 .26
BLEURT .13 .16 .22 .28 .20 .26 .22 .29 .19 .25
Table 4: (Step 1) Performance comparison of different models across different criteria. $\tau$: Kendall’s Tau, $\rho$: Spearman’s correlation, MSE: Mean Squared Error. Best results in each column are in bold. Higher is better for $\tau$ and $\rho$ and lower is better for MSE. The exact model versions can be found in the appendix. All values were computed over 200 samples and run to run variations were minimal. Three traditional machine translation baselines are given below.
@@
We first prompt the language model with the human-curated rubrics for each criterion and ask it to give a score. The LLM is also given one gold translation for reference. The full prompt can be found in Appendix D. As shown in Table 4, across all the criteria, LLM-based evaluation metric outperforms traditional metrics, however, even the best evaluator, GPT4-o does not match human-level correlation reported in Table 2. Our findings show that larger proprietary LLMs tend to do better than smaller models.
@@
In particular, a careful analysis reveals that GPT4-o and all other models completely fail to notice serious mistakes in honorifics. Table 3 shows that GPT4-o often gives high scores for translations that received low scores from humans regarding honorifics (notice the low F1 score for labels 1, 2 and 3).
@@
Korean honorifics play a crucial role in dialogue-based literary texts, drastically impacting the reading experience when improperly translated. Our evaluation rubric assigns a default score of 5 to non-dialogue passages, which means a significant number of passages automatically get 5. Thus, in our case, it is important for the models to detect when the honorifics get translated incorrectly. However, as Figure 6 shows GPT4-o’s predicted label distribution does not match well with human-to-human agreement. In contrast, GPT4-o does a better job at matching human preference distribution for Syntax, Content Accuracy, and Lexical choices (Figure 7), however, they still tend to provide overoptimistic estimations, a phenomenon captured by higher MSE compared to inter-human. We explore this phenomenon further in Section 6.
We conduct ablations of each component of the pipeline and reach the following conclusion:
  • CoT does not improve performance and, in fact, sometimes hurts (Table 9).
  • Specific grading rubric is important for some models, and not so much for others (Table 10).
  • Providing a reference translation does not help much and in fact, for some models, it is better not to provide one (Table 10).
  • Providing few shot examples has little to no effect (Table 11).
Appendix H digs into the observations in more detail.

### v3

@@ Abstract
In this work, we propose and test a two-step framework for evaluating literary machine translation, focusing on English to Korean. The framework combines RULER, a rubric-based evaluation tailored to language-specific features, with VERSE, a language-agnostic question–answering module. Our results show that it provides fine-grained, interpretable metrics suited for literary translation and achieves stronger correlation with human judgment than traditional machine translation metrics. However, it still falls short of inter-human agreement, particularly in culturally sensitive areas such as Korean honorifics. We also find that LLMs tend to favor translations produced by other LLMs. These findings highlight both the potential of our framework and the need for culturally aware evaluation methods to ensure accurate and sensitive translation of literary works. Code [^1] https://github.com/sheikhshafayat/ruler-verse and dataset [^2] https://huggingface.co/datasets/skshafayat/two-step-lit-eval are publicly available.
@@ 1 Introduction
While new approaches such as agentic systems are being utilized for literary translation Wu et al. (2024); He et al. (2024), evaluation remains challenging because automated metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020) are not capable of capturing the nuance of literary translation Thai et al. (2022).
Yet, for the agentic translation systems to work at scale, we need an automatic evaluation method first, which will give signal to the LLM agents and pipeline developer whether the translation meets the required criteria or not since human evaluation in each iteration is costly. Recently, however, pretrained LLMs have become more common in evaluating machine translation Kocmi and Federmann (2023); Freitag et al. (2023), in lieu of and in collaboration with human evaluation.
@@
Our analysis demonstrates that current LLMs approximate human preference in English–Korean literary translation on some dimensions, but they consistently underperform in culturally specific areas such as honorifics. This highlights both the potential of automated evaluation and its limits in capturing cultural features crucial for assessing literary translations.
@@ 2 Background
Building on James S. Holmes’s 1970s’ foundational theories of literary translation, scholars like Susan Bassett, André LeFevere, and Lawrence Venuti have led the “cultural turn” in translation studies, which emphasizes the importance of cultural and historical contexts in literary translation Snell-Hornby (2006). Literary translation has a special place within broader translation studies for its focus on the aesthetic and affective use of language as well as its engagement with fictional content. Some of the persisting concerns in literary translation include the tension between literal (word-for-word) and free or adaptive (sense-for-sense) Munday et al. (2022), the visibility of the translator Venuti (2017), and the social role of the translator Blakesley (2018). More recently, scholars including Kolb and Macken have examined the potential of the integration of technology and machine translation in literary translation processes Kolb (2023); Macken et al. (2022). Our work presents a pipeline incorporating the cultural and aesthetic aspects of literary translation into the evaluation of machine literary translation.
@@
String-based metrics like BLEU Papineni et al. (2002) have been widely used in machine translation evaluation. Similarly, more recent metrics such as BLEURT Sellam et al. (2020), COMET Rei et al. (2020), and XCOMET Colombo et al. (2023) leverage BERT-style encoder language models Devlin et al. (2019) fine-tuned on translation preference dataset. However, previous work has shown that these metrics are not suitable for literary translation Thai et al. (2022). Departing from earlier trends, Kocmi and Federmann (2023) employed the cutting-edge LLM GPT-4 OpenAI (2023) to assess machine translation, where their approach, GEMBA-MQM, set a new benchmark for translation quality estimation at WMT 2023 Freitag et al. (2023). Recent studies from Wu et al. (2024) and Chen et al. (2024) also use GPT-4 as a judge for literary translation for web novels and Chinese poetry, respectively.
We take a step back and conduct a comprehensive study on the LLMs’ capabilities of fine-grained evaluation for literary translation. A closely related work, Zhang et al. (2024), provide a broad human-study benchmark (LITEVAL-CORPUS) highlighting the inadequacy of MQM for literature and the limits of current automatic metrics.
@@
Our proposed question-answering-based translation evaluation method relates closely to automatic metrics like TIFA Hu et al. (2023), VPEval Cho et al. (2024), and DSG Cho et al. (2023) for measuring the consistency of text-to-image generation models such as DALL-E Ramesh et al. (2021) and Stable Diffusion Podell et al. (2024). In this line of work, an LLM first generates a list of verification questions from the given user’s prompt, for example “is there a car?” “is the car blue?” and then passes the questions to a vision language model such as BLIP Li et al. (2022) which answers the verification questions. This approach also has a resemblance to how previous works have measured faithfulness in summarization Kim et al. (2024); Wang et al. (2020) and long-form factuality evaluation Min et al. (2023). Another work has Krubiński et al. (2021) explored question-answering paradigms for machine translation evaluation before. In this work, we extend this paradigm to do fine-grained evaluation of literary translation.
@@ 3 Dataset, Metrics & Annotators
Human Annotators
Three literary experts with extensive experience in both English and Korean literature develop and validate the framework presented in Section 4. They also carry out the annotations in Section 5 to validate the framework.
@@ 4.1 Step 1: RULER
We first conduct a small-scale qualitative evaluation of machine-translated literary texts to identify the capabilities that evaluation models should possess, with a particular focus on the types of errors they should detect. For this, we select five stories (marked dark gray in Table 8) and translate them using various large and small language models, as well as commercially available translation services, like Google Translate, to get a diverse set of translations. We observe that LLMs show inconsistent quality in doing literary translation. Larger models like GPT-4o or Claude Sonnet are more likely to produce remarkably sophisticated translations, while smaller ones often make unexpected errors, even with relatively simple passages. Our analysis identifies four major types of errors machine translation systems make when translating literary works, some of which are especially pronounced in English-to-Korean translations: word/phrase/expression-level issues, context-related issues, issues relevant to honorifics, and form-related issues (see examples in Table 1). We provide a detailed explanation of these categories with concrete examples in Appendix D.
@@
The detailed rubric for each criterion will be released with our code and data after acceptance.
@@ 4.2 Step 2: VERSE
Eliminating common errors is insufficient for achieving high-quality literary translation. A successful translation must also maintain the essential and unique literary qualities of the original work, which can vary significantly between different texts. To capture such nuances, we propose VERSE, a question-answering-based evaluation framework. We prompt a high-performing LLM (in our case GPT-4o OpenAI (2024)) to generate a list of verification questions that must be satisfied in a good translation. We instruct the question-generating model to focus on the literary aspects of the passage, with a brief summary of the story provided for context. Figure 1 provides examples of the questions generated by the LLM.
@@
One expert human inspected 50 randomly sampled generated questions and found 44 of them to be “very helpful”. The six remaining questions, while relevant, were either too generic or unanswerable based on the provided context alone.
@@
To provide better interpretability for the LLM generated questions in VERSE, we classify the questions into nine categories using an LLM. For all our experiments, the questions are classified in a zero shot manner by GPT-4o. The categories, [^3] Notice that the categories are not necessarily mutually exclusive and one question might fit into more than one category. as shown in Figure 2, are defined by a human expert after qualitatively examining many evaluation questions. To validate the accuracy of the LLM classification, one human expert manually annotates 150 questions into the nine categories in which GPT-4o scores  77% Top-1 match in accuracy. An examination of the misclassified examples reveals that most misclassifications also fall into a closely related category (Figure 3).
5 Results: Do LLM judges catch common errors and literary nuance?
@@
We first create a diverse set of translations by machine translating the stories in our datasets with various LLMs and commercial translation service (e.g., Google Translate). The LLMs ranged from small 7B models to GPT-4o and Claude Sonnet 3.5. We also vary different prompting techniques — sentence level, paragraph level, providing extra summary, zero shot, and few shot — to generate diverse translation (see Appendix C). Human examination confirmed that the generated translation varied widely, ranging from barely fluent Korean with English code-switching behavior of small models to decent translations generated by large proprietary models.
We then randomly select 20 machine translated paragraph pairs from each of the 10 stories marked gray in Table 8 for human annotation (200 paragraphs in total). We also randomly select three evaluation questions corresponding to those paragraphs for annotation for Step 2 (600 questions in total).
Because of the subjective nature of evaluating literary translation, we create an extensive annotation guideline alongside the rubric so that annotators have the same understanding of the given task. We used the Label Studio platform [^4] https://labelstud.io/ for the annotation. The annotation guideline and annotation interface will be released with the codebase. Table 2 shows that the annotators have high inter-annotator agreement. We also ask the annotators to re-annotate a subset of 40 paragraphs after four weeks of the first annotation, and the intra-annotator agreement was  0.7.
RULER Kendall’s Spear. MSE Krip.
@@
Honorifics .70 .74 .86 .74
Lex. Choices .73 .81 .69 .80
Syntax .66 .76 .86 .75
Content .73 .81 .75 .80
VERSE .65 .71 .41 .70
Table 2: Inter-annotator agreement metrics across categories. Higher values indicate better agreement for Kendall’s $\tau$, Spearman’s $\rho$, and Krippendorff’s $\alpha$, while lower values are better for MSE. The values are averaged over three annotators. Notice that the last row, VERSE, belonging to Step 2, is measured between 1 to 3 while the top four rows, belonging to Step 1 (RULER), are measured on a 1 to 5 scale.
@@
We compare several open-weight and proprietary models (Table 9) as evaluators for our experiments. We set the temperature to 0.0 for all experiments in this section for reproducibility.
@@
According to the common practice of literary machine translation Karpinska and Iyyer (2023); Chen et al. (2024); He et al. (2024) , we report COMET Rei et al. (2022), BERTScore Zhang* et al. (2020), BLEURT Sellam et al. (2020) and GEMBA-MQM Kocmi and Federmann (2023) as the baseline. However, the comparison may not be entirely appropriate because these metrics’ designs are quite different from our framework, meant to be used in sentence-level evaluation, while our framework is designed for paragraph-level texts. See Appendix B.1 for details.
@@ 5.1 RULER: Can LLMs Capture Common Mistakes in Literary Machine Translation?
Honor. 1 30 0.79 0.07 0.81 0.67
2 4 0.15 0.10
3 9 0.11 0.04
4 9 0.22 0.14
5 147 0.91 0.88
Lexical 1 23 0.54 0.43 0.56 0.48
2 29 0.47 0.44
3 42 0.45 0.37
4 50 0.50 0.51
5 55 0.73 0.56
Syntax 1 24 0.61 0.39 0.52 0.44
2 32 0.55 0.44
3 47 0.50 0.27
4 45 0.37 0.43
5 50 0.62 0.58
Content 1 26 0.57 0.19 0.55 0.47
2 29 0.41 0.47
3 42 0.45 0.22
4 49 0.48 0.43
5 53 0.74 0.72
@@
Table 3: Performance comparison across different evaluation criteria. H-F1 represents the average F1-score between human annotator pairs (3 pairs), while LLM-F1 shows the average F1-score between human annotators and GPT-4o (3 pairs). H-Acc and LLM-Acc show the average accuracy for human-human and human-LLM comparisons respectively. Support (Sup.) shows the average number of instances for each label across three annotators (rounded).
Honorifics Lexical Syntax Content Step 2
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE
gpt-4o .51 .56 2.03 .68 .77 0.72 .62 .71 0.97 .70 .78 0.81 0.49 0.53 0.63
claude-3-5-sonnet .40 .41 2.22 .60 .67 1.13 .59 .66 1.17 .57 .64 1.35 0.38 0.41 0.63
claude-3-5-haiku .42 .45 1.80 .54 .59 1.35 .56 .64 1.22 .55 .62 1.39 0.39 0.42 0.66
gpt-4o-mini .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 0.37 0.40 0.63
claude-3-haiku .22 .23 2.17 .45 .50 1.55 .42 .47 1.47 .51 .56 1.51 0.19 0.20 0.87
gemma-2-27b-it .34 .38 2.08 .47 .52 1.27 .49 .56 1.24 .49 .55 1.25 0.18 0.19 0.70
gemma-2-9b-it .28 .30 2.36 .43 .48 1.40 .36 .41 1.47 .52 .58 1.34 0.09 0.09 0.73
llama-3-8b-it -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 0.05 0.06 1.06
COMET .10 .12 .28 .35 .25 .32 .28 .36 0.24 0.31
BERTScore .12 .15 .23 .29 .22 .28 .24 .30 0.20 0.27
BLEURT .14 .17 .23 .30 .21 .28 .23 .30 0.20 0.26
GEMBA-MQM (gpt-4o) .26 .32 .49 .63 .48 .61 .50 .63 0.38 0.46
Table 4: Evaluation results on RULER (Step 1) and VERSE (Step 2). Step 1 reports Kendall’s $\tau$, Spearman’s $\rho$, and mean squared error (MSE) across four fine-grained criteria: honorifics, lexical choice, syntax, and content. Step 2 provides overall correlation with human judgments of whether each criterion was satisfied. Higher is better for $\tau$ and $\rho$, and lower is better for MSE; Overall, larger models consistently outperform smaller ones, though performance varies across criteria and evaluation settings. Table 11 includes some newer models.
@@
We first prompt the language model with the human-curated rubrics for each criterion and ask it to give a score. The LLM is also given one gold translation for reference. As shown in Table 4, across all the criteria, our framework outperforms traditional metrics and GPT-4o based GEMBA-MQM, however, even the best evaluator, GPT-4o does not match human-level correlation reported in Table 2.
Honorific Failure Mode
@@
A careful analysis reveals that GPT-4o and all other models completely fail to notice serious mistakes in honorifics. Table 3 shows that GPT-4o often gives high scores for translations that received low scores from humans regarding honorifics (notice the low F1 score for labels 1, 2 and 3).
@@
Korean honorifics play a crucial role in dialogue-based literary texts, drastically impacting the reading experience when improperly translated. Our evaluation rubric assigns a default score of 5 to non-dialogue passages, which means a significant number of passages automatically get 5. Thus, in our case, it is important for the models to detect when the honorifics get translated incorrectly. However, as Figure 5 shows GPT-4o’s predicted label distribution does not match well with human-to-human agreement. In contrast, GPT-4o does a better job at matching human preference distribution for Syntax, Content Accuracy, and Lexical Choices. However, it still tends to provide overoptimistic estimations, captured by higher MSE compared to inter-human. We explore this phenomenon further in Section 6.